OpenStudy (anonymous):

What is the slope of the line 5x - 2y = 10? (4 points) none of the above

OpenStudy (anonymous):

Put the equation into the form y = mx + b and then "m" will be the slope.

OpenStudy (anonymous):

for the lnear eq of the form ax+by=c slope of the line is given by -(a/b) thus the slope hre is -5/(-2) =5/2